Give the domain, range, and zeros of the function f(x)=sqrt(9-x^2)
Would really appreciate any help, Thanks

Answer by lwsshak3(11628)

Give the domain, range, and zeros of the function f(x)=sqrt(9-x^2)
**
f(x)=sqrt(9-x^2)
Given equation is for the upper half of a circle with radius 3
domain:
radican≥0
9-x^2≥0
(3+x)(3-x)≥0
number line
<...-....-3....+....3....-.....>
Domain:[-3,3]
..
Range: [0,3]
..
Zeros:
9-x^2=0
x^2=±9
x=±√9=±3
x=3
or
x=-3